# Keyframer: Empowering Animation Design using Large Language Models

### v1

@@ 6.1. RQ2: What design strategies do users take to prompt LLMs for animations using natural language?
Collectively, our participants generated 205 unique prompts. [^12] Participants in total entered 221 prompts to the system. In some cases, the LLM would produce code with bugs and the participants were prompted to regenerate code using the same prompt. The count of unique prompts excludes those repetitive prompts. On average, each participant generated $15.815.8$ (SD = $4.04.0$) unique prompts during their user study session. Participant EP1 generated the most unique prompts ($2626$), and participant EP12 generated the least ($1111$). On average, the prompts consist of $16.716.7$ (SD = $15.215.2$) words, with the longest prompt consisting of $142142$ words and the shortest prompt consisting of only $22$ words (“remove exhaust”).
Among all the unique prompts, 51.2% are new prompts (i.e., prompts entered after clicking the “Add New Prompt” button), while 48.8% are edits to existing prompts (‘reprompts’). The majority of prompts (52.7%) requested refinement to the generated animation, while 42.9% requested new animation behavior for an item in the scene and 5.9% were a combination of the two. On average, each prompt asked for $2.02.0$ (SD = $1.61.6$) elements in the SVGs to be animated. The prompt that mentions the most number of elements contains $1010$ elements, while $33$ prompts do not mention any elements and only provide high level instructions such as “Generate a few animations for this scene.”
@@
We identified two different dimensions of prompting strategies: 1) decomposed versus holistic prompting, and 2) high specificity vs semantic prompting. Next, we describe the differences between these dimensions by illustrating with examples from our user study.
@@
Decomposed vs Holistic Prompting. We classified prompts as decomposed if they involved animating elements one by one through sequential prompts. In contrast, holistic prompts describe how multiple elements within a scene should be animated simultaneously within a single prompt.
All users took on a decomposed prompting strategy for at least one activity, alternating between A) prompts specifying new animation behavior for an item in the scene, and B) prompts requesting refinements to the generated animation. Figure 9 displays on the left how EP3 first requests new animation for the ‘specks’ (shown in green), followed by a refinement prompt to change the specks colors (in purple), then repeats for the Saturn ring. Through this process, users focused on animating a single feature and getting it to a state they were satisfied with before moving on to animating another element: “It’s how I’m used to doing animations—one piece at a time… I have to do this one animation, I have to tweak it, I have to do the next animation, and then I tweak it” (EP10).
@@
Holistic prompting was observed with 4 of the 13 users, with each user using this strategy during the second activity in particular  [^13] We believe holistic prompting may have been more popular in the second activity because the illustration lended itself to more sequenced-based animation compared to Activity 1, with users specifying how items in the scene should respond to the rocket lifting off.. With holistic prompting, users specified the behavior of multiple items simultaneously, often defining expected coordination or timing between elements in a single prompt (as shown on the right in Figure 9 where EP11 specified how the rocket, exhaust, and cloud should move in one prompt). Software engineer EP4 described this strategy by sharing, “It’s like sculpting. It was probably easier to try and get the rough shape of it all at the very beginning, and then fine tune the details versus trying to tell it cut by cut.” Participants EP1 and EP11 used holistic prompting because they believed it was necessary to have items to move simultaneously, rather than one after another. Participant EP10 also saw this strategy as a time saving step, describing how he would use the initial generated design to do further refinement in the properties editor.
@@
However, one challenge with holistic prompting was that in reprompting for multiple items to be animated at once, users run the risk of overriding desired behavior, as the temperature of the LLM requests was set to enable slight variations with the same prompt. This led users like EP5 to ask if it were possible to lock in certain parts of the code to prevent from being overwritten on regeneration.
In summary, the predominant use of decomposed prompting suggests that Keyframer enabled users to iteratively refine their designs through sequential prompting, rather than having to consider their entire design upfront. However, users that employed holistic prompting found a one-shot approach could effectively reduce time to generate by enabling more of the scene to be animated at once.
@@
High Specificity vs Semantic Prompting. We observed two different styles of describing animations: 1) high specificity and 2) semantic prompts. High specificity prompts, which represented 34.634.634.634.6% of all unique prompts, are characterized by using animation keywords (like opacity, rotate, and scale) as well as expected values (e.g., ‘between 0-1.5 seconds’). In contrast, the majority of prompts (84.4%) were semantic, which were less defined and more descriptive (e.g., ‘make the clouds wiggle’). [^14] We qualitatively coded the “high specificity” and “semantic” features among all 205 unique prompts. These features were not mutually exclusive, as some prompts could contain high specificity for one element and semantic descriptors for another; thus the percentages sum up over 100%. We present a taxonomy of semantic prompts types in Table 2, with prompts describing movement, timing, position, and coordination of objects in the scene, as well as incorporating analogies, indefinite values, and ideation requests, where users let the LLM come up with ideas (e.g., ‘make it look cool’). Figure 10 shows the breakdowns of high specificity and semantic prompts among the four user groups; we observed the greatest percentage of semantic prompts from the LCHA group, while the highest use of high specificity prompts comes from those with programming experience (HC).
Table 2. Taxonomy of Semantic Prompt Types
Prompt Type
Definition
Frequency
Examples
Movement
Describing translation
68.4%
@@
twinkle, wobble, balloon upwards
Timing
Describing speed
27.5%
@@
slowly, very fast, with an elastic feel
Position
Describing relative placement
30.1%
@@
from below, vertically on the same plane
Coordination
Relationships between items
19.7%
together, keep attached
Other Visual
@@
Opacity, color change, fonts
8.8%
increasingly transparent, alternate color
Analogies
Similes or metaphors
2.6%
like stars in the sky, like an active fire
Indefinite Values
Indeterminate values
14.0%
at different rates, random value
Ideas
Requesting new ideas from the LLM
11.9%
do something crazy, show alternatives
Figure 10. Percentages of high specificity and semantic prompts by participant groups.
[…]
Participants were surprised by how well the LLM could interpret semantic prompts. Participant EP7, who used words like ‘morph,’ ‘shimmer,’ and ‘grow’ in his prompts, stated, “I was generally impressed by how ambiguous I could be with my descriptions of the animation. It usually picked the correct transform to use, or a reasonable transform to use.” This idea was echoed by EP8, who used the prompt Give me 3 designs where the clouds wiggle: “The fact that it got make the clouds wiggle in an interesting way, in a way that I was happy with, is kind of wild.” This was surprising especially to users who questioned whether they needed to use specific animation keywords: “I would be curious if I can use human language, like pulse or twinkle or if I have to use, like, turn and rotation—more technical words” (EP9).
Others shared how being less specific opened up creative possibilities, potentially increasing the likelihood the LLM would return unexpected results. Participant EP13 described how he intentionally avoided using standard animation terminology in his prompts, believing that doing so might “give it [the LLM] more room to try something creative.” In another example, when EP5 did not have a specific design in mind, he would prompt the system with a general instruction (such as “Make it look cool”) to curate new ideas: “I am not that sure what to do sometimes… instead of explaining the steps of making it look as cool as possible at an atomic level…instead of being all imperative about it, you can just say what the goal is…[the LLM] can just fill in the rest.” Participant EP4 summarized: “If you could be more vague [in your prompts], then you have more options for surprise and delight.”
For others, semantic prompts were a time saving step, reducing the need to spend time defining all parameters upfront. For example, participant EP8, who transitioned from high specificity to more semantic prompts during her session as shown in Figure 11, said, “It’s a lot easier to just assume it [the LLM] will do a pretty good duration, or that I could mess with it easily, so I don’t need to type it all the way through.”
Overall, we observed how semantic prompts enabled users to focus on higher level goals, both because the LLM had reasonable interpretations of their requests and because it returned sensible defaults that users could directly refine through Keyframer’s editors.
Figure 11. High specificity prompts (such as the one on the left) use animation keywords and values such as opacity and duration. Low specificity prompts (such as the two on the right) are more descriptive and more open to interpretation. These three prompts were all used by participant EP8, who started off with high specificity prompts but then transitioned to using semantic prompts.
[…]
@@
Debugging Prompts. While participants were frequently surprised that Keyframer understood their prompts, they expressed how articulating motion in natural language was challenging in and of itself. There were several moments where the LLM output did not match a user’s expectations, yet they acknowledged that the LLM had a reasonable interpretation: “It’s definitely that sort of language barrier thing, where it did exactly what I wanted, but it turns out that wasn’t what I wanted. It did exactly what I told it to do, but I should have told it to do something else” (EP4).
When the output did not match their expectations, it could be unclear if the source of the problem was the prompt or the LLM’s capabilities: “I’m not sure if it was me or the program” (EP3). In these cases, the one sentence explanation generated alongside each code snippet, along with semantically descriptive generated keyframe names and animation properties, could aid users with verifying whether their prompts were interpreted correctly: “I like that it’s writing it out. I can see what it’s doing or how it’s taking my commands” (EP9). Seeing the streaming output could also give users ideas for how to prompt in the future: “As it’s generating code, you can kind of look into the black box of what is actually happening. If you see a field [name], that can be useful… if I want to change that property, that’s the word I can use” (EP5). Additionally, the ability to generate multiple designs unexpectedly helped participant EP4 debug his prompt ‘Generate 3 images of the rocket flying over the moon.’ When he found that none of the generated designs matched his vision, he shared: “I clearly need to find a way to get it [the LLM] to understand what the moon is. With the 3 different outputs, it definitely ignored me, so I need to hone in on getting it to recognize that.” Asking the LLM to generate three designs thus enabled him to confirm across multiple examples the need to refine his prompt.
Several users wanted the LLM to confirm its understanding before generating designs. For example, EP5 suggested a ‘more interactive process’: “It [the LLM] could give me feedback like, ‘Do you mean this? Or ‘Do you want to do this instead?”’ Others proposed having the system make suggestions during the prompting process: “If I type in ‘moon,’ highlight the moon [in the image preview], but also in the text itself, make the word moon be colored so that I understand that it’s a keyword” (EP4). Providing this type of visual feedback may help users gain more confidence that the system is capable of understanding keywords and requests.
@@ 6.2. RQ3: How does Keyframer support iteration in animation design?
In the following sections, we focus on how Keyframer shapes users’ iteration process—including exploration and refinement—on their animation designs.
6.2.1. Exploration
Participants regard Keyframer as useful in helping them brainstorm different ideas for animations, rating the system with a $3.73.7$ (SD = $0.60.6$) on average, which is between ‘helpful’ (4) and ‘neutral’ (3).
@@ 6.2.1. Exploration
Surprise LLM output facilitates new ideas. At times, Keyframer surprised participants with a design that they had never thought of. Participants tended to value such surprises as they could spur new ideas. For example, EP8 once prompted the LLM to generate a design in Activity 2 where the clouds wiggle, whereas the LLM generated animations to the rocket as well. Although it was different from what EP8 has asked for, she ended up liking the animation and built subsequent designs on top of it. The generated output itself can seed future prompting; participant EP1, who is new to animations, shared, “I’m reacting to what I see as its output and then adjusting on my end…in some ways, it’s informing and giving me ideas for what to explore next.” Similarly, EP3 summarized her experience with the unexpected designs Keyframer generated: “Even if it doesn’t come out the way you want it to, it’s showing you another path that you can take… I think it’s helpful to just come up with some other creative solutions.” For those more experienced in designing animations, Keyframer could offer them new directions for technical implementation of animation effects, potentially broadening their animation skills. For instance, when EP7 prompted the LLM to create animation on the flame of the rocket as the rocket launches, the LLM generated code that stretched the flame vertically, which was not what he had in mind: “I thought I’d have to morph the flame to get something like this effect… I was like, ‘Oh, this isn’t bad either.’ It’s basically like another person suggesting something for me to try.”
@@
LLMs for generating design variations. The other approach was to explicitly ask the LLM to generate multiple designs, such as “generate 3 designs where the cloud wiggle.” A total of 9 out of the $1313$ participants prompted the LLM to generate multiple designs. Participants thought the option to have the LLM generate multiple designs at once was helpful for brainstorming and comparing ideas: “It was nice to kind of see the different possibilities and then feel like I could choose among them which one was more closely aligned with” (EP1).
The option to generate multiple designs could be effective in different stages of the animation design process. In the very beginning of the design process where participants were exploring what to animate, they would leverage it to map different directions. As EP4 explained, “I know roughly what I want, but then if I could get it [the LLM] to initially show me like a larger set of possibilities, that would probably help me change the vision of what I want, because it’s showing possibilities and then maybe more surprises.” In later stages of the process, when participants were making decisions on certain details or parameters of the animation, they also valued experimenting with variations. For example, when EP2 was deciding on the speed of the animation of certain elements, he asked the LLM to generate variants to compare and contrast the effect of different duration values. In another example, EP5 thought generating design variants was a good way to overcome “artist block” when they encountered “friction” in the design process, and this could happen at the beginning of the process when they did not know where to start and towards the end when they were unsure what to do next.
Tensions with generating variants. Although some participants found the option to prompt the LLM to generate multiple designs helpful, they could be hesitant to use the feature in practice. When participants already had a clear idea of what they wanted, they would not generate multiple designs as it might steer them away from what they envisioned. Participants also reflected on their existing workflow when approaching creative solutions, where they typically focus on improving one element at a time instead of going broad. Participants from an engineering background, in particular, found a tension between the option to explore multiple design variations and the option to keep building on a single design: “I think it’s probably part of being an engineer and trained to do things iteratively and see what happens, particularly with a non familiar system…it’s a troubleshooting thing. If I ask it to do 10 things, I don’t even know if it’s going to do one correctly first” (EP4). Participant EP8 further explained the different needs they have when engineering versus designing animations: “When I’m thinking in the more design side, you don’t know what you want until you see it sometimes. So it really helps to have a couple variations on what it looks like… If it’s more pure engineering, I probably would just like the one design.” Participants also wished that the variants could better match their designs goals, be more distinct, and take less time to generate. For example, EP5 shared, “I was hoping that it [the LLM] would maybe be willing to [take] more liberties with some of these suggestions,” which aligned with feedback from EP2 and EP7 that sometimes only 1 or 2 of their variants were viable. We did not see any users request more than 3 designs and believe this may be because generating multiple designs leads to longer response times.
6.2.2. Refinement
Participants in general consider Keyframer a helpful tool in assisting them with refining animations, rating $3.83.8$ (SD = $1.01.0$) on average, between “helpful” (4) and “neutral” (3). We observed two ways that the participants refine their animation designs using Keyframer: through editing code and editing prompts.
@@ 6.2.2. Refinement
Code editing interfaces offer granular creative control. Our system supports participants with all levels of programming experience with making edits to LLM-generated code, even for those without programming background (i.e., those in the LCHA and LCLA groups). A majority of participants (69.2%) edited the LLM-generated code with either of Keyframer’s editors. Seven participants edited the code directly using the Code Editor feature, while $77$ edited the code using the Properties Editor feature. Five participants used both editors, and $44$ participants did not make any edits to the code generated by LLM. On average, each participant made edits to 25.125.125.125.1% (SD = $0.20.2$) of the instances where they received code generated by the LLM. Table 3 shows the editing actions by participant group. Notably, 4 out of 6 Low Code users edited output code, and of the 5 High Code users who edited LLM output, the majority (80.0%) used the Properties editor, suggesting that the Properties editor can be beneficial to those with all levels of programming experience. In fact, most participants that used the editors used both the Code and Properties editor throughout their use of Keyframer (55.6%), suggesting that each can be beneficial regardless of prior experience with animation or code.
Participants preferred to make direct edits on the code generated by the LLM when they had clearly defined goals: “It was nice that I was able to then go into the code and edit things that I didn’t want to ask the model to do because I already knew exactly what I wanted” (EP4). For participants that were less familiar with CSS like EP2, the Properties Editor gave them an easier option to make these edits: “I thought it did a really good job at parsing the code and then giving me kind of a WYSIWYG type editor for the code. I think that’s really helpful for people that aren’t comfortable jumping into the code.”
Table 3. Characteristics of code edits by participant group
Participant % of users who used % of users who used % of users who did not Mean % of code instances
group the Code editor the Property editor edit code at all with user edits
LCLA 33.3% 66.7% 33.3% 16.7%
HCLA 33.3% 33.3% 66.7% 13.8%
HCHA 75% 75% 0% 35.6%
LCHA 66.7% 33.3% 33.3% 17.4%
Combination of reprompting and code editing fuels design iteration. For those who chose to not edit the output code such as EP1 and EP4, editing prompts was an alternative way to modify the animation designs. Those who did not edit code output, like EP4, still acknowledged the potential benefit of making those edits, but chose not to due to the novelty effect of using a prompt-based interface and the time limit of the activity: “The prompt is the novelty, whereas hand editing [the code], there’s no magic in that…I can absolutely do that by hand, but it’s not as fun.”
During the sessions, we observed the common practice of both re-prompting and editing output to refine animation designs. Participants shared the strategy of first using natural language to prompt for a design close to what they had imagined, and then leveraging the editing interfaces to adjust details: “I feel like the natural language part gets me 95% of the way there, and then the last 5% I would prefer to do edits here [the Code Editor]” (EP9).
Some valued using the LLM to generate boilerplate code that they could build off. CSS experts like EP7 appreciated that LLM-generated code saved them from having to look up syntax documentation: “Remembering all these names [of CSS properties] and writing them out always is annoying…it’s nice that it [the LLM] shows all the transform origins and iteration count so I didn’t have to think about that.” Building off boilerplate code gives creators more time to focus on the design itself: “You can save so much time using a tool like this versus typing everything out. It allows you to do more stuff because you have more time to iterate” (EP13).
In fact, participants agreed that it was the combined powers of prompt input and code editing that made Keyframer helpful for refinement. For example, EP13 shared how he found himself largely refining his animations through manual code edits, rather than via prompting because the output editors reduced the amount of time to refine his animations.

### v2

@@ 6.1. RQ2: What design strategies do users take to prompt LLMs for animations using natural language?
Each participant entered on average $15.8$ prompts (SD = $4.0$, min = 11, max = 26), with a mean word length of $16.7$ words per prompt (SD = $15.2$). The longest prompt consisted of $142$ words and the shortest contained only $2$ words (“remove exhaust”).
@@
In this section, we describe two different dimensions of prompting strategies we identified: 1) decomposed versus holistic prompting, and 2) high specificity vs semantic prompting.
6.1.1. Decomposed vs Holistic Prompting
@@
. We classified prompts as decomposed if they involved animating elements one by one through sequential prompts. In contrast, holistic prompts describe how multiple elements within a scene should be animated simultaneously within a single prompt.
All users took on a decomposed prompting strategy for at least one activity, alternating between prompting for new animation behavior for an element and requesting refinements to the generated animation, as shown on the left of Figure 8. Through this process, users focused on a single feature and getting it to a state they were satisfied with before moving on to another element. This strategy enabled users to iteratively refine their animations through sequential prompting, rather than having to consider their entire design upfront in a single prompt. Keyframer’s ability to support decomposed prompting workflow is particularly effective for animation design as it enables designers to preview outputs, refine the movement of specific elements, and then proceed to the next step. Further, the generated output from an earlier prompt could seed future prompting; participant EP1, who is new to animations, shared, “I’m reacting to what I see as its output and then adjusting on my end…it’s informing and giving me ideas for what to explore next.” Decomposed prompting enabled users to adaptively refine their design goals as they built up an animated scene through individual prompts.
@@
In contrast, holistic prompting was observed with 4 of the 13 users, with each user using this strategy during the second activity in particular [^16] We believe holistic prompting may have been more popular in the second activity because the illustration lent itself to sequenced-based animation, e.g., specifying how items in the scene should respond to the rocket lifting off.. With holistic prompting, users specified the behavior of multiple items simultaneously in a single prompt, as shown on the right in Figure 8. Software engineer EP4 described this strategy by comparing it to sculpting: “It was probably easier to try and get the rough shape of it all at the very beginning [in a single prompt], and then fine tune the details versus trying to tell it [the LLM] cut by cut.” The other 3 participants used holistic prompting to save time or because they felt it was necessary for having items move simultaneously. However, by revising the same underlying prompt, users run the risk of overriding desired behavior.
@@
6.1.2. High Specificity vs Semantic Prompting
@@
. We observed two different styles of describing animations. High specificity prompts (34.634.6% of all prompts) are characterized by using animation keywords (like opacity, rotate, and scale) as well as expected values (e.g., ‘between 0-1.5 seconds’). In contrast, the majority of prompts (84.4%) were semantic, which were less defined and more descriptive (e.g., ‘make the clouds wiggle’). [^17] We qualitatively coded the “high specificity” and “semantic” features among the 205 unique prompts used across all users. These features were not mutually exclusive, as some prompts could contain high specificity for one element and semantic descriptors for another; thus the percentages sum up over 100%. We present a categorization of semantic prompt types in Table 2, with categories 1) describing the movement, timing, direction, coordination, or other visual qualities of objects in the scene, and 2) the phrasing of those requests through incorporating analogies, using indefinite values, or asking for ideas (e.g., ‘make it look cool’).
Table 2. Categorization of Semantic Prompt Types
Prompt Type Definition Frequency Examples
@@
Movement Describing transformation (translation, rotation, scaling, etc.) 68.4% twinkle, wobble, balloon upwards
@@
Timing Describing speed and timing curves 27.5% slowly, very fast, with an elastic feel
@@
Direction Describing relative placement and orientation 30.1% from below, vertically on the same plane
Coordination Relationship and synchronization between items 19.7% together, keep attached, follow
@@
Other Visual Opacity, color change, fonts 8.8% increasingly transparent, alternate color
Analogies Similes or metaphors 2.6% like stars in the sky, like an active fire
Indefinite Values Indeterminate values 14.0% at different rates, random value
Ideas Requesting new ideas from the LLM 11.9% do something crazy, show alternatives
Participants were surprised by how well the LLM could interpret semantic prompts: “I was generally impressed by how ambiguous I could be with my descriptions. It usually picked the correct or reasonable transform to use” (EP7). Similarly, EP8, who used the prompt Give me 3 designs where the clouds wiggle shared: “The fact that it got make the clouds wiggle in an interesting way that I was happy with is kind of wild.” Being less specific opened up creative possibilities: “If you could be more vague [in your prompts], then you have more options for surprise and delight” (EP4). For example, EP5 utilized open-ended prompts (e.g.,“Make it look cool”) to curate new ideas: “Instead of explaining the steps of making it look as cool as possible at an atomic level…you can just say what the goal is and [the LLM] can just fill in the rest.” For others, semantic prompts saved time, reducing the need to define all parameters upfront. Participant EP8 transitioned from high specificity prompts in Activity 1 (e.g., For each of the specks, animate them pulsing between 85% and 115% of their current size. Each pulse should take between .5 and 1.5 seconds”) to more semantic prompts by Activity 2 (e.g., Move the rocket exhaust and clouds from the bottom to the top), sharing how generated properties like duration were often reasonable as is or were easy enough to directly edit, thus not requiring full specification upfront.
6.1.3. Debugging Prompts
@@
. While participants were frequently surprised that Keyframer understood their prompts, they expressed how articulating motion in natural language was challenging in and of itself. There were several moments where the LLM output did not match a user’s expectations, yet they acknowledged that the LLM had a reasonable interpretation: “It’s definitely that sort of language barrier thing where it did exactly what I wanted, but it turns out that wasn’t what I wanted. It did exactly what I told it to do, but I should have told it to do something else” (EP4). Several users wanted the LLM to confirm its understanding before generating designs by asking the user to select one of multiple interpretations. Others proposed having the system make suggestions during the prompting process by highlighting corresponding keywords from the prompt alongside its corresponding SVG element as they typed, indicating whether the LLM could correctly link words in their prompt to items in the scene.
@@ 6.2. RQ3: How does Keyframer support iteration in animation design?
In the following sections, we focus on how Keyframer shapes users’ iterative process—including exploration and refinement—on their animation designs.
6.2.1. Exploration through unexpected output and generating variants
@@
Participants found Keyframer useful for helping them brainstorm directions for animations, both because of unexpected design it generated along with the use of the multiple variants feature. At times, Keyframer surprised participants with a design they had never thought of, which spurred new ideas: “Even if it doesn’t come out the way you want it to, it’s showing you another path that you can take… I think it’s helpful to just come up with some other creative solutions” (EP3). Keyframer could also offer new directions for technical implementation. For instance, when EP7 prompted the LLM to animate the flame of the rocket as it launches, the system returned a vertically scaling effect that surprised him: “I thought I’d have to morph the flame to get something like this effect…I was like, ‘Oh, this isn’t bad either.’ It’s basically like another person suggesting something for me to try.”
@@
A total of 9 out of the $13$ participants also explored ideas by explicitly ask the LLM to generate multiple designs. This design feature helped users brainstorm and compare ideas: “It was nice to see the different possibilities and then feel like I could choose among them which one was more closely aligned” (EP1).
Generating multiple designs could effectively support different stages of the animation design process. When participants were initially exploring what to animate, they could leverage the multiples feature to map different directions: “If I could get it [the LLM] to initially show me a larger set of possibilities, that would probably help me change the vision of what I want” (EP4). In later stages, participants valued using variations to support detailed refinement like timing and duration of animated elements.
At the same time, some participants expressed hesitancy about generating multiple designs in practice. When participants already had a clear idea of what they wanted, they would not generate multiple designs as it might steer them away from their vision. Some wished that the variants could better match their designs goals, be more visually distinct from one another, and take less time to generate, with EP2, EP5, and EP7 expressing that sometimes only 1 or 2 of the generated variants were relevant to their request.
6.2.2. Refinement through combined prompting and direct editing
@@
Keyframer’s editors supports participants with a range of programming experience with editing LLM-generated code. A majority of participants (69.2%) edited the LLM-generated code with either of Keyframer’s editors. Seven participants edited code using the Code Editor, while $7$ made edits with the Properties Editor. On average, each participant edited 25.125.1% (SD = $0.2$) of the instances where they received generated code. Notably, 4 out of 6 Low Code users edited output code, and of the 5 High Code users who edited LLM output, the majority (80.0%) used the Properties editor, suggesting that the Properties editor can be beneficial to both experts and novices. In fact, most participants that used the editors used both the Code and Properties editor throughout their use of Keyframer (55.6%), suggesting that each editor can be beneficial regardless of prior experience with animation or code.
Participants described how it was the combined powers of prompting and code editing that made Keyframer helpful for refinement. A common strategy was to first use natural language to prompt for a design close to what a user had imagined and then leverage the editing interfaces to adjust details: “I feel like the natural language part gets me 95% of the way there, and then the last 5% I would prefer to do edits here [the Code Editor]” (EP9). Participants also shared that getting boilerplate code generated from natural language could leave them more time to focus and iterate on the design itself. Especially when having clearly defined goals, participants preferred to edit the generated code directly compared to re-prompting the system.